In a school with 500 students, the average height is 165 +-10 cm. Assuming that the height follows the normal distribution:

If we choose a sample of 25 people from the above school:

What is the probability that a student is taller than 169 cm?

What is the probability that a student is less than 163 cm tall?


1
Expert's answer
2021-05-07T11:51:02-0400

Solution:

Given, "\mu=165,\sigma=10,n=25"

"X\sim N(\mu,\sigma)"

(a) "P(X>169)=P(z>\dfrac{169-165}{10/\sqrt{25}})=P(z>2)=1-P(z\le2)"

"=1-0.97725=0.02275"

(b) "P(X<163)=P(z<\dfrac{163-165}{10/\sqrt{25}})=P(z<-1)=0.15866"